On an object of mass 1kg moving along x-axis with constant speed 8m/s a constant force 2N is applied in positive y-direction.Find its speed after 4 seconds.?

Grade:12th pass

2 Answers

Debdeep
6 years ago
Velocity along x=8(constant) Acceleration along y=2m/s^2Velocity along y after 4 s=2*4=8m/sSpeed =8\/2m/s
Deligant
5 years ago
  1. Mass=1kg.     Velocity= 8i^.   Force=2j^.  F= ma  so.   2j^=1.v-u/t.  = v-8i^/4.  So 8j^+8i^=v. Since he is asking the magnitude.    root over 82+82=8root 2